OpenStudy (anonymous):

How do you solve (x - 2y)(x + 2y) = 4

OpenStudy (anonymous):

good question

OpenStudy (bahrom7893):

No solutions

OpenStudy (anonymous):

multiply out, then subtract 4, then factor or use quadratic formula

OpenStudy (bahrom7893):

2 unknowns one equation.. not enough info i should say..

OpenStudy (anonymous):

oh woops!

OpenStudy (anonymous):

couldn't hurt to read the problem. what is there to solve? one equation, two unknowns.

OpenStudy (anonymous):

Base off of this information (x - 2y)(x + 2y) = 4 which quanity is greater A) x^2-4y^2 or B)8

OpenStudy (bahrom7893):

8

OpenStudy (anonymous):

\[(x^2-4y^2=4\] \[y^2=\frac{4-x^2}{4}\] \[y=\pm\frac{\sqrt{4-x^2}}{2}\]

OpenStudy (anonymous):

ok now i am getting cranky so maybe i should log off. really though before wasting peoples time it would really be nice if the entire question would be posted.

OpenStudy (bahrom7893):

Good point satellite

OpenStudy (anonymous):

that is good way to solve it, thanks and the answer is 8